OpenStudy (anonymous):

For classes 0f 278 students, find the mean and standard deviation for the number born on the 4th of july. ignore leap years.

OpenStudy (preetha):

Hi Irish

OpenStudy (anonymous):

Hi

OpenStudy (fibonaccichick666):

Hey there, we are unable to assist until you post the full question. Please try not to leave out any details :)

OpenStudy (preetha):

Do you understand what std dev is?

OpenStudy (preetha):

And Fib is right. Do you have the data table?

OpenStudy (anonymous):

I have the data table

OpenStudy (preetha):

Tell me what is in the data table, what are the column headings? Or post a pic here.

OpenStudy (michele_laino):

we need to know the distribution of the students

OpenStudy (anonymous):

That is all I have for the question, I'm sure the table I have is correct but I will attach it

OpenStudy (michele_laino):

that's the binomial distribtuion

OpenStudy (preetha):

I am not sure you have the question represented here correctly. What does it say in your question in your book? Or screenshot it.

OpenStudy (fibonaccichick666):

Is this perchance a multiple part question, above or in a previous question is there a description of the students?

OpenStudy (preetha):

Have to agree with Fibonacci again. Irish, we want to help you. Fib and Michele are two of our best helpers.

OpenStudy (anonymous):

how do I do a screen shot so you can see it

OpenStudy (preetha):

Command+shift+4 for macs.

OpenStudy (preetha):

Or use your phone.

OpenStudy (preetha):

Or just copy and paste the question

OpenStudy (anonymous):

I have to find the value of mean and the valuee of standard deviation of this problem

OpenStudy (preetha):

Right. To calculate the mean of a set of values, you add all the numbers together and divide this sum by the number of values. SO if you have 100 data points, or values, and their sum is 2340. The mean is 2340/100.

OpenStudy (preetha):

So you will have a table with the values. Do you see it?

OpenStudy (fibonaccichick666):

or you will have a question saying something like, the class has this many students with birthdays in july this many in december, on this day, etc.

OpenStudy (anonymous):

no i dont have the table with the values

OpenStudy (michele_laino):

hint: we can consider these probabilities: one student is born on 4th of July with a probability of 1/365 and a student is not born on 4th og July with the probability of 364/365 so we have p= 1/365, q = 364/365 and mean value is \[\mu = N \times p{\text{ = 278/365 = }}...{\text{?}}\] whereas the standard deviation is: \[\sigma {\text{ = }}\sqrt {Npq} = \sqrt {278 \times \left( {1/365} \right) \times \left( {364/365} \right)} = ...?\]

OpenStudy (anonymous):

that is the equation I used but I can't figure out the whole problem. The answer I got was 0.7595571

OpenStudy (michele_laino):

please wait, I'm checking the answers...

OpenStudy (fibonaccichick666):

Since the mean is used to calculate standard deviation, you will need to make sure your answer is accurate there first.

OpenStudy (michele_laino):

I got: \[\mu = N \times p{\text{ = 278/365 = 0}}{\text{.7616}}\] which value is very close to yours. Your problem is a simple application of the binomial distribution. Please you have to keep in mind the meaning of the binomial distribution

OpenStudy (preetha):

Irish, did you understand what Michele did?

OpenStudy (anonymous):

sort of

OpenStudy (preetha):

Which part is troubling you?

OpenStudy (anonymous):

if her answer is the mean how do you get the standard deviation

OpenStudy (fibonaccichick666):

Can you reread the section Michele has titled hint? Do you understand where those formulas come from?

OpenStudy (michele_laino):

the computation, in order to get the formula of the standard deviation, starting from the equation of probability, is not easy. I think that it is more important to understand the problem, rather than the algebra which conduct us to the formula for the standard deviation.

OpenStudy (anonymous):

I'm just going to my textbook and read more about it to understand it cuz I'm really confused now. Thanks for all your help

OpenStudy (michele_laino):

for example, if I pick a student from your sample of 278 students, then I have a probability p=1/365 that student was born on 4th of July, and a probability of 364/365, that student was not born on 4th of July.

OpenStudy (michele_laino):

Now, using some of combinatorial calculus, we can show, that probability distribution, is given by the binnomial distribution

OpenStudy (michele_laino):

oops..binomial*

OpenStudy (fibonaccichick666):

I think that may be above irish's head. @Irish69 if I may make a suggestion, take 10 minutes, get a snack, come back after you have cleared your head, and between @Michele_Laino Preetha and I, we will be able to help you understand.

OpenStudy (michele_laino):

please if you need more help, send me a message! @Irish69

OpenStudy (preetha):

Thanks all for your help. @Michele_Laino and @FibonacciChick666

TheSmartOne (thesmartone):

Here is a tutorial on how to take screenshots if you need to in the future :) http://openstudy.com/study#/updates/54e3dca9e4b0f2455c58e57f